how do you find the product of (9x-7)(9x+4)
Answer by tommyt3rd(5050)

Distribution is the property that allows us to multiply polynomials

(9x-7)(9x+4)=
9x(9x+4)-7(9x+4)=
81x^2+36x-63x-28=
81x^2-27x-28=


:)